If a+b is a root of X square + ax + b equals 0 then what is the maximum value of b square?
1
Expert's answer
2019-05-20T02:46:09-0400

If a+b is a root of x^2+ax+b=0, then

"(a+b)^2+a(a+b)+b=0"

Hence,

"a^2+2ab+b^2+a^2+ab+b=0"

which yields to

"b^2+b(3a+1)+2a^2=0"

Therefore,

"a^2+6a+1\geq 0"

and

"b_{1}=\frac{-(3a+1)+\sqrt{a^2+6a+1}}{2},""b_{2}=\frac{-(3a+1)-\sqrt{a^2+6a+1}}{2}"

We have

"a\in(-\infty,-3-2\sqrt{2})\cup(-3+2\sqrt{2},+\infty)"

If

"a\in(-\infty,-3-2\sqrt{2}),"

then b_{1}^{2}>b_{2}^{2}. Hence, in this case b_{1}^{2} is the maximum value of b^2. Otherwise, if

"a\in(-3+2\sqrt{2},+\infty),"

then b_{2}^{2}>b_{1}^{2}. In this case, b_{2}^{2} is the maximum value of b^2.